OpenStudy (anonymous):

$9500 is invested, part of it at 12%, and part of it at 9%, For a certain year, the total yield is $1032.00. How much was invested at each rate? How much was invested at 12% How much was invested at 9%

OpenStudy (anonymous):

System of Equations: a + b = 9500 .12a + .09b = 1032 From the first equation: a = 9500 - b Plug that value of 'a' into the second equation. .12(9500 - b) + .09b = 1032 Distribute on the left side: 1140 - .12b + .09b = 1032

OpenStudy (anonymous):

Combine like terms: 108 = .03b Solve for b through division of .03: b = 3600.

OpenStudy (anonymous):

Take that value for 'b' and plug it into the FIRST equation: a + 3600 = 9500 Solve for 'a' by subtracting both sides by 3600: a = 5900. a) $5900 at 12% b) $3600 at 9%

OpenStudy (anonymous):

Don't forget to select a Best Response. Thanks and take care.